# iBERT: Interpretable Style Embeddings via Sense Decomposition

### v1

@@ A.1 Broader Applications
Rather than relying on keyword overlap or dense similarity, queries can be modified in latent space to match stylistic attributes (e.g., formality, emoji use). This enables retrieval that is both semantically relevant and stylistically aligned—particularly useful in assistant and RAG pipelines where tone coherence matters.
@@ A.4 Training and Implementation Details
All iBERT models use k=8k{=}8 senses, 768-dim hidden states, and FlashAttention-2 (Dao, 2024) for fast attention computation. AdamW optimizer with learning rate of ($2\!\times\!10^{-5}$), a batch size of 99 is used. The models have an early stopping with 3-epoch patience over validation loss and Flash-Attention-2 is used for faster compute. We perform both masked language modeling and sentence embedding training at input lengths of 128 and 512 tokens. MLM uses 5% of FineWeb (750B tokens), and contrastive fine-tuning is performed using anchor-positive-negative triplets from StyleDistance (SD) and Wegmann (WG) datasets.
@@
All models are initialized with encoder weights from ModernBERT (Warner et al., 2025), replacing into either a Vanilla (standard) BERT or our proposed iBERT architecture. Due to the larger Context Sense Block in iBERT, we freeze the encoder weights in an initial warmup stage (for both baseline and iBERT to maintain parity) and allow the heads to converge. Once stabilized and converged, we unfreeze the full model and resume end-to-end training until convergence. This process is used in both MLM and contrastive phases. The details of the MLM stage are listed in Table 8.
@@ A.4.1 Sentence Embedding Training and Replications
For sentence-level contrastive training, we train a total of five distinct sentence encoder models (per data configuration): a baseline SBERT built on top of Vanilla-BERT, and four sentence encoders built on iBERT-MLM checkpoints using different pooling frameworks. These correspond to mean-pooling (v1, $\tau{=}\infty$), top-sense (v2, $\tau{=}0$), softmax-blend (v3-1, $\tau{=}1$), and peaked-blend (v3-10, $\tau{=}10$). Each is treated as a separate model, not as a parameter sweep. All models are trained at two input lengths (128 and 512) and evaluated on StyleDistance and Wegmann datasets.
To account for training variability, we perform 5 independent training replications for each sentence encoder (Vanilla-SBERT and all iBERT variants) at both token lengths (128 and 512). Reported results reflect mean and standard deviation across the runs.
Training runs $=5$ $\left\{\begin{array}[]{@{}l@{}}\text{model types:}\\ \text{Vanilla, iBERT-v1, v2}\\ \text{iBERT-v3-1, v3-10}\end{array}\right$.
@@ A.5 Style Contrast Visualization
We visualize the separation of contrastive style pairs (e.g., number substitution, personal pronoun frequency) using t-SNE projections. For each example, we compare Vanilla-SBERT embeddings (left) and iBERT-v2 embeddings (right). In both cases, sentence embeddings were obtained after training on StyleDistance.
@@
A.6 Non-Target Style Ablation Visualization
A.7 Global t-SNE for Style Clustering
(a) Vanilla-SBERT (128 tokens)
@@ A.7 Global t-SNE for Style Clustering
Figure 6: t-SNE projections of sentence embeddings for all 40 style contrast pairs. Each point represents a sentence, color-coded by its style label. Vanilla-SBERT shows more dispersed and overlapping clusters, while iBERT-v3-10 shows tighter, aligned groupings — indicating improved style modularity and disentanglement. Both these models were trained on StyleDistance triplets.
We visualize sentence embeddings across all 40 style contrasts using t-SNE, comparing Vanilla-SBERT and iBERT-v3 in Figure 6. The iBERT model exhibits sharper geometric structure, with more linear and segregated style clusters. This supports our hypothesis that modular pooling (v3) induces interpretable and axis-aligned representations, in contrast to the entangled embedding space of Vanilla-SBERT.
A.8 Measuring Locality of Sense-Based Editing
For a given style $x$ (e.g., Text Emojis / No Emojis), we identify the sense $\ell^{*}$ with the highest probing activation and ablate it during encoding by zeroing the corresponding sense gain. Let $P_{x}$ denote the positive examples for style $x$, and $P_{x}^{\text{orig}}$, $P_{x}^{\text{edit}}$ be their embeddings before and after ablating $\ell^{*}$, respectively. We compute cosine distance between a set of embeddings $A$ and a centroid $\mu$ as:
$d(A,\mu)=\frac{1}{|A|}\sum_{a\in A}\left(1-\cos(a,\mu)\right).$ (1)
The target editability (reported in Table 3) is then:
$\Delta_{x}=\frac{d(P_{x}^{\text{orig}},\mu_{x})-d(P_{x}^{\text{edit}},\mu_{x})}{d(P_{x}^{\text{orig}},\mu_{x})},$ (2)
where $\mu_{x}$ is the centroid of negative examples for style $x$.
To quantify whether sense ablation incurs unwanted shifts on other style dimensions, we repeat this computation for every $y\neq x$ (with centroid $\mu_{y}$):
$\Delta_{y}=\frac{d(P_{x}^{\text{orig}},\mu_{y})-d(P_{x}^{\text{edit}},\mu_{y})}{d(P_{x}^{\text{orig}},\mu_{y})}.$ (3)
We report the following aggregate locality metrics:
Avg. Other Shift $=\frac{1}{|\{y\neq x\}|}\sum_{y\neq x}\Delta_{y},$ (4)
Max Other Shift $=\max_{y\neq x}\Delta_{y}.$
Intuitively, a useful and disentangled iBERT sense should yield a large, positive $\Delta_{x}$ while keeping Avg. Other Shift and Max Other Shift close to zero, indicating minimal collateral influence on unrelated style axes.

### v2

@@ A.1 Broader Applications
Rather than relying on keyword overlap or dense similarity, queries can be modified in latent space to match stylistic attributes (e.g., formality). This enables retrieval that is both semantically relevant and stylistically aligned—particularly useful in assistant and RAG pipelines where tone coherence matters.
@@ A.2 Effect of Aggregation Sharpness
Variant Frozen Phase Unfrozen Phase Total
Time Loss / Epochs Time Loss / Epochs Time Epochs
128 iBERT 23h 1.6912 / 57 3d 6h 1.3365 / 85 4d 5h 142
BERT 2h 1.3607 / 5 1d 2h 1.2228 / 31 1d 5h 36
512 iBERT 2d 9h 1.4418 / 46 1d 7h 1.2941 / 16 3d 16h 62
BERT 14h 1.1876 / 6 3d 8h 1.0940 / 46 3d 22h 52
Table 8: Training runtimes, epochs, and final losses for iBERT-MLM and BERT (MLM) under 128- and 512-token limits. Epoch counts exclude early-stopping patience rounds.
@@ A.4 Training and Implementation Details
All iBERT models use k=8k{=}8 senses, 768-dim hidden states, and FlashAttention-2 (Dao, 2024) for fast attention computation. AdamW optimizer with learning rate of ($2\!\times\!10^{-5}$), a batch size of 99 is used. The models have an early stopping with 3-epoch patience over validation loss and Flash-Attention-2 is used for faster compute. We perform both masked language modeling and sentence embedding training at input lengths of 128 and 512 tokens. MLM uses 5% of FineWeb (750B tokens), and contrastive fine-tuning is performed using anchor-positive-negative triplets from StyleSynth / StyleDistance (SD) and Wegmann (WG) datasets.
@@
All models are initialized with encoder weights from ModernBERT (Warner et al., 2025), replacing into either BERT or our proposed iBERT architecture. Due to the larger Context Sense Block in iBERT, we freeze the encoder weights in an initial warmup stage (for both baseline and iBERT to maintain parity) and allow the heads to converge. Once stabilized and converged, we unfreeze the full model and resume end-to-end training until convergence. This process is used in both MLM and contrastive phases. The details of the MLM stage are listed in Table 8.
@@
Datasets, Licensing, and Safety
Our study is based on publicly available datasets, which have been used in prior research on author attribution and stylistic analysis. These datasets originate from online platforms and may inherently contain offensive or personally identifiable content. However, we do not apply any additional filtering beyond what has been done in previous studies. To ensure ethical usage, we follow the guidelines set by the original sources that published these datasets and acknowledge any potential biases or content-related concerns that may arise.
Additionally, we discuss the licensing terms associated with the datasets used in our experiments. The datasets are provided under open-access licenses, allowing their use for academic research. We ensure compliance with these licenses and properly attribute the sources in our work. Detailed information on dataset licensing is included in the final version of the paper.
@@ A.4.1 Sentence Embedding Training and Replications
For sentence-level contrastive training, we train a total of five distinct sentence encoder models (per data configuration): a baseline SBERT built on top of BERT, and four sentence encoders built on iBERT-MLM checkpoints using different pooling frameworks. These correspond to mean-pooling (v1, $\tau{=}\infty$), top-sense (v2, $\tau{=}0$), softmax-blend (v3-1, $\tau{=}1$), and peaked-blend (v3-10, $\tau{=}10$). Each is treated as a separate model, not as a parameter sweep. All models are trained at two input lengths (128 and 512) and evaluated on StyleSynth and Wegmann datasets.
To account for training variability, we perform 5 independent training replications for each sentence encoder: (S)BERT and all iBERT variants, at both token lengths (128 and 512). Reported results reflect mean and standard deviation across the runs.
Training runs $=5$ $\left\{\begin{array}[]{@{}l@{}}\text{model types:}\\ \text{(S)BERT, iBERT-v1, v2}\\ \text{iBERT-v3-1, v3-10}\end{array}\right$.
@@ A.5 Style Contrast Visualization
We visualize the separation of contrastive style pairs (e.g., number substitution, personal pronoun frequency) using t-SNE projections. For each example, we compare (S)BERT embeddings (left) and iBERT-v2 embeddings (right). In both cases, sentence embeddings were obtained after training on StyleSynth.
@@
A.6 Global t-SNE for Embedding Clustering
(a) BERT (128 tokens)
@@ A.6 Global t-SNE for Embedding Clustering
Figure 6: t-SNE projections of sentence embeddings for all 40 style contrast pairs. Each point represents a sentence, color-coded by its style label. BERT shows more dispersed and overlapping clusters, while iBERT-v3-10 shows tighter, aligned groupings — indicating improved style modularity and disentanglement. Both these language models were first trained on FineWeb and then trained on StyleSynth triplets.
We visualize sentence embeddings across all 40 style contrasts using t-SNE, comparing BERT and iBERT-v3 in Figure 6. The iBERT model exhibits sharper geometric structure, with more linear and segregated style clusters. This supports our hypothesis that modular pooling (v3) induces interpretable and axis-aligned representations, in contrast to the entangled embedding space of BERT.
A.7 Model Schematic and Embeddings
We conclude the appendix with Figures 5 and 6, which illustrate the full model architecture and the global embedding structure.